If the interior angle of a regular polygon is 5 times the exterior angle, then how many sides does the polygon have? Please tell me how to do it.
Answer by stanbon(57407)

If the interior angle of a regular polygon is 5 times the exterior angle, then how many sides does the polygon have? Please tell me how to do it.
------------
An interior, exterior pair are always supplementary.
You have x + 5x = 180
6x = 180
x = 30
-------
The exterior angle is 30 degrees.
--------------------------------------------
Note: The sum of the exterior angles is always 360 degrees.
--------------------------------------------
# of exterior angles = 360/30 = 12
------------------
# of sides = # of exterior angles = 12
==================================
Cheers,
Stan H.